OpenStudy (judson_rockets):

What is the surface area of this rectangular prism? A. 836 in2 B. 520 in2 C. 209 in2 D. 418 in2

OpenStudy (abdullahm):

Find the area of each rectangle and add them all up. There are two rectangles with a width of 5 and a height of 8. There are two rectangles with a width of 5 and a height of 13. There are two rectangles with a width of 8 and a height of 13. Find the area of all the rectangles and add them up. :)

OpenStudy (judson_rockets):

209

OpenStudy (abdullahm):

That's wrong. :(

OpenStudy (abdullahm):

What is 5 * 8 * 2 = ? We're multiplying it by 2 since there are two rectangles like that. What is 5 * 13 * 2 = ? What is 8 * 13 * 2 = ? Add up those three numbers to get the correct final answer. :)

OpenStudy (judson_rockets):

80 For the first one 130For the second one and 208 For the third one

OpenStudy (judson_rockets):

So it would be 418

OpenStudy (abdullahm):

That is correct!